OpenStudy (anonymous):

What is the product in simplest form? State any restrictions on the variable.

OpenStudy (mikeyymonroe):

??? product of what?

OpenStudy (anonymous):

\[\frac{ 4a ^{5} }{ 7b ^{4} }\times \frac{ 2b ^{2} }{ 2a ^{4} }\]

OpenStudy (mikeyymonroe):

hmm lemme see if i can do this on a seperate pice of paper brb.

OpenStudy (anonymous):

okie doke thanks xo

OpenStudy (mikeyymonroe):

hmm this problem is kinda iffy with me at the moment... do you mind if I ask for a picture?

OpenStudy (mikeyymonroe):

ok thank you i will try and solve it ASAP

OpenStudy (anonymous):

okay thanks

OpenStudy (mikeyymonroe):

my best guess is A becuase it is multiplication so you add the powers. but dont go immidiately with my answer

OpenStudy (anonymous):

okay thank you!

OpenStudy (mikeyymonroe):

of course anything to help.

OpenStudy (mikeyymonroe):

thank you for the medal :) very kind of you